110 joules of heat is added to a gaseous system, whose internal energy is 40j; then the amount of external work done is


        (a)    150 J


        (b)    70 J


        (c)    110 J


        (d)    40 J


Grade:10

1 Answers

Simran Bhatia
10 years ago

(b)

DQ = DU + DW

Þ DW = DQ - DU = 110 - 40 = 70 J